OpenStudy (anonymous):

2^x=3^x-1 please show steps

OpenStudy (regan):

3^(x-1) or (3^x)-1 ?

OpenStudy (anonymous):

3^(x-1)

OpenStudy (anonymous):

srry

OpenStudy (regan):

ok that makes this alot easier : ) log(2^x)=log(3^(x-1)) loging both sides x*log(2)=(x-1)log(3) log laws x=(x-1)*log(3)/log(2) i will call log(3)/log(2)=k x=k(x-1) x=kx-k 0=kx-k-x k=kx-x k=x(k-1) x=k/(k-1) where k=log(3)/log(2)